OpenStudy (anonymous):

What are the relationship among angle measure in degrees, angle measure in radians, and arc length?

random231 (random231):

360 degrees = 2*pi so, 1 degree= (2*pi)/360

OpenStudy (anonymous):

you should go w/ random231

random231 (random231):

 arc length AB = x*pi*radius